OpenStudy (anonymous):

Factorise completely 18c + 24d

OpenStudy (australopithecus):

what is a common multiple of 18 and 24. what number can multiple with one number to give 18 and another number to give 24

OpenStudy (australopithecus):

when I say what number I mean what one number.

OpenStudy (anonymous):

What you can maximum do in this is to take 6 common out of the two terms: 6(3c + 4d)

OpenStudy (australopithecus):

can multiply with two different numbers

OpenStudy (anonymous):

I am sorry for my behaviour waterineyes

OpenStudy (anonymous):

On the last question

OpenStudy (anonymous):

It is okay..

OpenStudy (cwrw238):

18 = 2 * 3 * 3 24 = 2 * 3 * 4 greatest common factor = 2 * 3 = 6 so take 6 outside the brackets to give 6(3c + 4d)

OpenStudy (anonymous):

Answer to australaphotecus is 6

OpenStudy (anonymous):

That is a small little girl just drawing a painting of me... @Messi10

OpenStudy (anonymous):

cwrw238 what do you mean

OpenStudy (australopithecus):

right so 6(3c + 4d) = 18c + 24d the factorized form is 6(3c + 4d) it wouldn't make sense to factor out any variables as both terms in the equation contain different variables

OpenStudy (anonymous):

Thanks